OpenStudy (anonymous):

Find the LARGEST of three consecutive integers such that 3 times the sum of the first and the third integer is equal to 30 more than 4 times the second integer. A) no such integers exist B) -28 C) 13 D) 16

OpenStudy (tkhunny):

Give it a name!

OpenStudy (anonymous):

what?

OpenStudy (tkhunny):

It's a simple suggestion. This guy right here: "LARGEST of three consecutive integers " Give it a name.

OpenStudy (anonymous):

i'll act like that made sense?

OpenStudy (anonymous):

the answer's 16.

OpenStudy (tkhunny):

It does make sense. You've probably done it many times. Why is this time harder? Just give it a name. We need to call it something, since "LARGEST of three consecutive integers " is a bit laborious.

OpenStudy (tkhunny):

Did you pull that out of a hat?

OpenStudy (anonymous):

no it's the answer.

OpenStudy (midhun.madhu1987):

Let the numbers be x , x+1 and x + 2. According to question, 3(x+x+2) = 4(x+1) + 30 3(2x+2) = 4(x+1) + 30 Solve this to find x. And the largest number will be x+2

OpenStudy (anonymous):

i'm doing it online, i already answered the question. the answer was 16.

OpenStudy (tkhunny):

No, let the greatest number by x and the other two x-1 and x-2.

OpenStudy (midhun.madhu1987):

okay... just an added information.. :)

OpenStudy (midhun.madhu1987):

all will be the same only....

OpenStudy (anonymous):

@tkhunny i don't know what you're saying it gave me the correct answer.

OpenStudy (midhun.madhu1987):

Whether it is x,x+1 and x+2 OR x, x-1 or x-2 they are all the same..

OpenStudy (tkhunny):

You can also call them Steve, Ralph, and Bob, but that's quite a bit less useful.

OpenStudy (tkhunny):

A correct answer is good, but knowing what you are doing is even better. You showed no evidence here that you know. You just pulled 16 out of the air.

OpenStudy (anonymous):

i clearly don't know how to do it but you're not making any sense.

OpenStudy (midhun.madhu1987):

 Can u prove that this is wrong?

OpenStudy (anonymous):

don't really care. have a great night. :)

OpenStudy (tkhunny):

I just can't tell what isn't making sense. You need to find a number, so call it something. x = The Number I want. That's all it takes.

OpenStudy (anonymous):

what i am wondering is why the insistence on algebra i would guess and check 10+11 + 12 is not big enough, i bet i would get it on the second try

OpenStudy (tkhunny):

"Guess and Check" is officially NOT algebra?

OpenStudy (midhun.madhu1987):

Every human has their own thinking ability .. that is what i have to say.. There is no specific way to do problems in algebra..

OpenStudy (tkhunny):

@midhun.madhu1987 I have no objection to your methodology. Just trying to make sure we don't all get trapped in just one way of doing things.

OpenStudy (midhun.madhu1987):

yes.. me too agreee to that.. i also suggested the same... Thanks..